OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Describe how the -3 was calculated and what it means.

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To }\mathbb{Open}\textbf{Study}}\] Best of luck for ur journey here in openstudy

OpenStudy (anonymous):

common difference =-3 Sum of AS = (first term + last term)(no. of terms) /2 first term =4 last term =-11 number of terms=x 4-(-11) = (-3)x x=5 therefore, there is 4 minus 3 five times to reach -11 som there is six terms Sum of AS =(4-11)(6)/2 =-21

OpenStudy (anonymous):

Thanks